how to solve this using substitution and check
x-6y+4=0
y= -3

Thanks

Answer by stanbon(75887)

solve this using substitution and check
x-6y+4=0
y= -3
-----
Solve for "x":
x - 6*-3 + 4 = 0
x +12 + 4 = 0
x = -16
===============
Cheers,
Stan H.
============